# II. Carbon-enhanced metal-poor starsthanks: All M3DIS 3D RHD CEMP models are available under https://nlte.mpia.de

### v1

@@ 2.3 Spectrum synthesis
Figure 2: Comparison of mean temperature stratification between the DISPATCH 3D CEMP model and the 1D MARCS model on the optical depth scale. Left: $\rm T_{eff}=5750\ K,\ log(g)=4.5$, Right: $\rm T_{eff}=5250\ K,\ log(g)=3.0$. Horizontal bars at the bottom correspond to the approximate formation region of the G-band in 1D and 3D.
Synthetic spectra in 1D and 3D are computed using the updated version of MULTI3D code (we refer to Eitner et al. 2024 for details). The code includes extensive tables of bound-free and free-free radiative cross-sections for 92 chemical elements and partition functions from Irwin (1981). The occupation formalism is adopted for H I, following the HBOP routine by Barklem & Piskunov (2003). The same code is used for the calculations with the MARCS and with M3DIS model atmospheres, to ensure consistency between results.
We synthesize the CH G-band in LTE between $\rm 4297\ and\ 4303\ \AA$ with a wavelength spacing of $0.001\ \AA$ . In 3D, we downsample the full cube horizontally to a resolution of $\rm 20\times 20$ points using linear interpolation. Such a horizontal down-sampling is sufficient to provide reliable profiles of spectral lines with abundance errors within $\sim\rm 0.01\ dex$ compared to simulations with a much higher horizontal resolution (Bergemann et al. 2019, 2021; Rodríguez Díaz et al. 2024; Eitner et al. 2024). The vertical resolution is not downsampled in the spectrum synthesis, but rather upsampled to an adaptive grid in the optically thin layers, to better resolve regions with large opacity gradients. For more information see Hoppe et al. (in prep.). The final, re-sampled models have a resolution of $20\times 20\times 180$ points.
@@ 3 Results
First we compare the scaled-solar 3D RHD and 1D HE model atmospheres, computed using identical stellar parameters ($\rm T_{eff}$, $\rm\log g$, metallicity). For simplicity, and to highlight the main physical differences, we discuss spatially averaged (on planes of constant optical depth) temperature-density (T-$\rm\rho$) profiles in this section but we emphasize that all subsequent spectrum synthesis calculations are performed in full 3D using 3D cubes. In Fig. 1 we show the T-$\rho$ (top) and T-$\tau$ (bottom) structures, where $\tau$ is the optical depth, of the MS (left, panels A and C) and of the sub-giant (right, panels B and D). As a representative case, we show both models at $\rm[Fe/H]=-5$ to investigate their properties and to illustrate the typical behaviour observed across the full set of models. All panels illustrate the respective atmosphere between 4-4 and 0 in $\rm\log\tau_{500}$.
@@
Comparing the 3D structures with their 1D counterparts, the dominant differences are apparent towards the upper layers of the atmospheres. For the MS model (Fig. 1, panels A and C), the mean 3D temperature structure starts to deviate from that of the 1D HE model around optical depth $\rm\log\tau_{500}\sim-2$. The $\rm 3D-1D$ difference reaches values of $\rm\sim-1\,500\ K$ at $\rm\log\tau_{500}=-4$. In those layers, the 3D RHD model is significantly and systematically cooler than the 1D MARCS model, which is well-known and was extensively discussed in the earlier work on 3D models of very metal-poor (VMP) MS and red giant stars (Gallagher et al. 2016, 2017; Collet et al. 2018). This difference is due to the fundamental property of convection, a process that is inherent to 3D RHD models, but is missing or parameterized by the mixing-length theory (Gustafsson et al. 2008) in 1D HE models. In metal-poor stars, the atmospheric structure is dominated by adiabatic cooling of rising, and hence expanding, fluid parcels of the gas. This is in stark contrast to stars with solar-like metallicity: in such conditions, the high metallicity leads to significant radiative heating (Stein & Nordlund 1998), which counteracts the adiabatic cooling effect and implies that the average 3D structure remains nearly in radiative equilibrium (e.g. Nordlund et al. 2009; Freytag et al. 2012; Magic et al. 2013a, b).
@@
A very similar behaviour can be seen in the structure of the 3D model of the sub-giant star (Fig. 1, panels B and D). The 3D RHD model is over $\rm\sim-1\,000\ K$ cooler compared to its MARCS counterpart in the layers with $\log\tau_{500}\lesssim-3$. Although the effect is very large, it is slightly less compared to that of the MS 3D model in the uppermost layers ($\log\tau_{500}\sim-4$). One would naively expect that a sub-giant 3D model, due to lower pressure, would show large deviations from the 1D HE model, compared to the MS case. However, there is another interesting effect. The lower density and cooler structure of this model favours formation of different molecules, including CH, OH, and CO. Due to radiative re-heating, the temperature balance of the 3D model hence is slightly closer to that of 1D HE MARCS model. We emphasize that this difference is not a direct consequence of the C-enhancement, it is also present in models with scaled-solar composition, but of the overall absence of metal absorbers in the energy budget of the atmosphere.
3.1.2 3D CEMP vs. 3D scaled-solar
@@
Comparing C-enhanced and scaled-solar 3D models directly (Fig. 1, blue crosses and blue, solid lines, respectively), we find that for the VMP main-sequence star (panels A and C), the effect is modest compared to the $\rm 3D-1D$ differences described above. At optical depth of $\log\tau_{500}>-3$, there is no significant difference, while towards the upper boundary of the atmosphere, the mean temperature of the CEMP model is $\rm\sim 400\ K$ hotter at $\rm\log\tau_{500}\sim-4$. In the sub-giant model (panels B and D), the same effect of C-enhancement is noticeable already at $\rm\log\tau_{500}\sim-2$, with temperature differences of $\rm\sim 200-300\ K$ throughout the entire optically thin part of the atmosphere. Overall, the effect on the main sequence model is $\rm\sim 100\ K$ larger at around $\rm log\ \tau_{500}\lesssim-4$ than on the sub-giant. Towards deeper layers, however, $\rm\log\tau_{500}\gtrsim-3$, the scaled-solar sub-giant remains consistently cooler than its CEMP counterpart, while the departures between the MS models are already negligible.
@@
The reason for this additional heating in the CEMP models is found in the increased abundance of C, which implies increased CH, CO and CN molecular abundances, a higher molecular line opacity, and consequently more radiative heating of the upper atmospheric layers. Because O – as well as N – remain at their scaled-solar values, specifically the CH G-band is one of the main contributors to the line opacity beyond $\rm\log\tau_{500}\sim-2$. In cooler atmospheres with very low [Fe/H], this effect is most noticeable. First of all, this is because already in the scaled-solar approximation C is an order of magnitude more abundant than Fe. In addition, the CH number densities are higher in cooler systems, as the chemical equilibrium favours a higher $\rm CH/C$ for lower T values. The re-heating of the upper atmospheric layers induced by the C-enhancement hence produces the opposite of the $\rm 3D-1D$ effect described in Sec. 3.1.1. We furthermore note that there is no such effect in 1D HE MARCS models. To highlight this, we include one example CEMP MARCS model with the same stellar parameters in panels A and C.
@@
Because of its lower $\rm log(g)$, the sub-giant model overall has a lower density at the same temperature as the main sequence model. A consequence of this is that the formation height of spectral lines generally shifts towards deeper layers in terms of optical depth. This explains the larger departures ($\rm 3D\ CEMP-\text{scaled-solar}$ models) at optical depth $\rm\log\tau_{500}\gtrsim-3$ among the sub-giant models compared to the main sequence models (comparing the difference between blue lines and blue crosses in panel C with panel D) through the shift of the molecular line heating contribution to deeper atmospheric layers. Generally, it may be noted that decreasing the overall effective temperature causes the opposite effect, as the increased molecular number densities increase the opacity and hence shift the contribution function to lower optical depth (but see also the discussion in appendix A).
@@
Figure 3: Difference between 3D CEMP and 1D MARCS models as a function of metallicity. Top: Relative temperature difference (percent). Middle: Logarithmic density differences (dex). Bottom: CH/H number density fraction, where $\rm n^{*}=n_{CH}/n_{H}$ is the CH number density relative to the hydrogen number density. Blue, dashed lines correspond to $\rm T_{eff}=5750\ K,\ log(g)=4.5$ models, red, solid curves to $\rm T_{eff}=5250\ K,\ log(g)=3.0$. The differences in each figure are computed at the representative formation height of the strong lines in the G-band. See Sec. 3.2 for more information.
@@
Figure 4: CH G-Band computed with M3DIS 3D and 1D MARCS model atmospheres. Black solid lines represent the CEMP 3D model, blue dashed lines the corresponding scaled-solar 3D model. Red solid lines show the 1D MARCS model with the same stellar parameters. Top: $\rm T_{eff}=5750\ K,\ log(g)=4.5$, bottom: $\rm T_{eff}=5250\ K,\ log(g)=3.0$.
3.1.3 3D CEMP vs. 1D HE
@@
Combining the results from Sec. 3.1.1 and Sec. 3.1.2, we now compare our M3DIS 3D CEMP model atmospheres with 1D HE scaled-solar MARCS models. The difference between these two model types is particularly important, as many studies rely on 1D scaled-solar MARCS models for spectral synthesis and abundance analysis. This raises the question of whether this simplifying assumption introduces systematic effects in the derived abundances.
@@
To find the answer, we first show the full 3D temperature structure of both selected 3D CEMP models at $\rm[Fe/H]=-5$ in Fig. 2. For both models, MS (left-hand side) and sub-giant (right-hand side), we used the C-enhancement of $\rm[C/Fe]=+3$. The 1D MARCS models are shown with a red line. The full 3D structure is indicated by the blue colour shading. We furthermore show in the inset a horizontal slice of the same atmospheres, interpolated to the optical surface at $\rm\log\tau_{500}=0$. We note the very different scales and granular size between the MS (left panel) and the sub-giant (right panel) models. The model of the MS star is, because of much larger gravity, about 20 times smaller compared to that of a sub-giant simulation, the latter exceeding 100 Mm in size. As described in the previous sections, the difference between 3D CEMP and 1D MARCS models is most striking in the optically thin layers of the atmosphere, where 3D CEMP models are still significantly cooler, even though the difference is reduced by the C-enhancement-induced re-heating. Both models converge towards the optical surface, the agreement is best in proximity of $\rm\tau=2/3$, indicating a very similar effective temperature of 1D and 3D models.
@@
Towards the bottom of Fig. 2 we furthermore include horizontal bars, indicating the formation region of the G-band. For this, we select a representative CH line, compute the contribution function in its core for each model atmosphere, and use the location of its peak as proxy for the optical depth at which the G-band typically forms. We draw the horizontal lines until $\rm\tau_{500}=1$, which corresponds to the formation of weaker features, line wings and the continuum. We emphasize that the computation of the G-band formation region does not affect the results or conclusions of this study; it is intended solely as a useful reference to aid in the analysis of model differences and to support their physical interpretation. In Fig. 2, it may be noted that the G-band forms slightly deeper within the atmosphere of the sub-giant as compared to the MS model, which is a consequence of the significantly lower surface gravity and hence overall lower density. This shift happens despite the lower effective temperature of the 3D sub-giant model. In the 1D model the upper layers are strongly dominated by radiative equilibrium – and hence the effective temperature – which means the drop in temperature between the MS and sub-giant models is more significant, the opacity rises and the formation region of the G-band shifts in the opposite direction.
@@
The evolution of the differences between 3D CEMP and 1D MARCS models with metallicity can be investigated in Fig. 3. For clarity, we refrain from showing the differences between the full profiles as a function of optical depth for each metallicity. Instead, we present a representative 3D–1D difference for each metallicity, computed from the atmospheric regions most relevant to G-band formation. Accordingly, the data shown in this figure is evaluated at the G-band formation height in each respective atmosphere, as described in the previous paragraph.
@@
In panel A of Fig. 3 the relative temperature difference is shown for the sub-giant models (red, solid lines) and for the MS models (blue, dashed lines). We find that 3D CEMP models are on average between $20\%$ and $25\%$ cooler than 1D HE models. At metallicities 6-6, 5-5, and 4-4, there is no significant variation in the 3D–1D difference. At $\rm[Fe/H]=-3$ the differences start to decline, and reach values around $5\%$ at metallicity 2-2. Besides forming in cooler parts of the atmosphere, the G-band furthermore forms in regions of lower density (panel B), with $\rm\log\rho_{3D}/\rho_{1D}$ between 0.6-0.6 dex at $\rm[Fe/H]=-6$ and 0.3-0.3 dex at $\rm[Fe/H]=-2$. Note that this is true despite the fact that metal-poor 1D HE models generally are less dense at the same optical depth (due to the hotter temperatures) and the data in panel B is mostly reflecting the differences in formation height through the exponential growth of density with depth.
@@
Additionally, it may be noted that the MS models show 3D CEMP effects similar to – if not even slightly larger than – the sub-giant models, especially when considering the temperature differences (Fig. 3, panel A). However, at this stage it is important to consider that the molecular equilibrium, which ultimately determines the CH number densities and hence directly influences the strength of the spectral lines in the G-band, is not a linear function of temperature and density, but rather exhibits a temperature and density dependence as dictated by the Saha equation (e.g. Carson 1992)
@@
where $\rm n_{x}$ is the number density of species $\rm x$, $\rm z_{x}$ the partition functions and $\rm D_{CH}$ the dissociation energy of the CH molecule, which is set to $\rm 3.47\ eV$ (Popa et al. 2023; Storm et al. 2025, and references therein). To highlight this non-linear dependence, we include panel C showing the fraction of CH/H number densities $\rm n^{*}_{3D}/n^{*}_{1D}$ – where $\rm n^{*}=n_{CH}/n_{H}$ – in the same formation region as panels A and B. Here it becomes obvious, that the number density of CH/H molecules in the sub-giants deviate by a factor of $\rm\sim 1.2$ more between 3D and 1D than what is found in MS models. So despite showing a slightly smaller 3D effect in temperature and density, the 3D effect is more prominent in the CH/H number density of sub-giants, which will ultimately be reflected in the G-band and $\rm[C/Fe]$ abundance corrections (see Sec. 4).
@@
Overall, it may be noticed that with increasing [Fe/H] the differences in the structure between 3D CEMP and 1D HE models decrease noticeably. The reason is that with increasing metal content, radiative heating starts to dominate the overall energy budget of the photosphere. The consequence is that adiabatic cooling becomes less dominant in the outer layers, which ultimately closes the gap between the average 3D structure and 1D HE MARCS models, which do not include this cooling mechanism in the first place.
@@
Our results are in a good agreement with work presented by Gallagher et al. (2016, 2017), although a different code (CO5BOLD, Freytag et al. 2012) was used in that study and their choice of stellar parameters does not fully overlap with that of our models. Gallagher et al. (2017) furthermore note that if the O abundance is changed, this has an impact on the number densities of the CO molecule. Consequently, the amount of free C available to form CH is decreased. Whereas generally it would be of interest to test non-solar abundances of different elements, in this work we do not explore the dependence on the $\rm C/O$ ratio. First, because there is evidence for CEMP stars with only very modest O enhancement (Bessell et al. 2004; Iwamoto et al. 2005). Also, most diagnostic OH features that are suitable for O abundance measurements (Asplund & García Pérez 2001) are located in the UV range, which is extremely sensitive to NLTE effects in metal-poor stars (Bergemann & Nordlander 2014; Lind & Amarsi 2024). Furthermore, no 3D NLTE $\rm C/O$ ratios are available for CEMP stars (Collet et al. 2018). We also note that Gallagher et al. (2017) focus on a hot sub-giant with metallicity 3-3, which leads to smaller differences between 1D and 3D models – partly because of the higher metallicity, and partly because of the lower $\rm CH/C$ ratio in both CEMP and non-CEMP models. We will explore the impact of different metallicities and stellar parameters on synthetic stellar spectra and C abundance determinations in the following sections.
3.2 Spectrum Synthesis of the CH G-band

### v2

@@ 2.3 Spectrum synthesis
Figure 2: Comparison of the mean temperature stratification between the DISPATCH 3D CEMP model and the 1D MARCS model on the optical depth scale. Left: $\rm T_{eff}=5750\ K,\ log(g)=4.5$. Right: $\rm T_{eff}=5250\ K,\ log(g)=3.0$. Horizontal bars at the bottom correspond to the approximate formation region of the G band in 1D and 3D.
Synthetic spectra in 1D and 3D are computed using the updated version of MULTI3D code (we refer to Eitner et al. 2024 for details). The code includes extensive tables of bound-free and free-free radiative cross-sections for 92 chemical elements and partition functions from Irwin (1981). The occupation formalism is adopted for H I, following the HBOP routine by Barklem & Piskunov (2003). The same code is used for the calculations with the MARCS and with M3DIS model atmospheres, to ensure the results are consistent.
We synthesised the CH G band in LTE between $\rm 4297\ and\ 4303\ \AA$ with a wavelength spacing of $0.001\ \AA$ . In 3D, we down-sampled the full cube horizontally to a resolution of $\rm 20\times 20$ points using linear interpolation. Such a horizontal down-sampling is sufficient to provide reliable profiles of spectral lines with abundance errors within $\sim\rm 0.01\ dex$ compared to simulations with a much higher horizontal resolution (Bergemann et al. 2019, 2021; Rodríguez Díaz et al. 2024; Eitner et al. 2024). The vertical resolution is not down-sampled in the spectrum synthesis, but rather up-sampled to an adaptive grid in the optically thin layers, to better resolve regions with large opacity gradients. For more information, see Hoppe et al. (in prep.). The final, re-sampled models have a resolution of $20\times 20\times 180$ points.
@@ 3 Results
First we compared the scaled-solar 3D RHD and 1D HE model atmospheres, computed using identical stellar parameters ($\rm T_{eff}$, $\rm\log g$, and metallicity). For simplicity, and to highlight the main physical differences, we discuss spatially averaged (on planes of constant optical depth) temperature-density (T-$\rm\rho$) profiles in this section but we emphasise that all subsequent spectrum synthesis calculations are performed in full 3D using 3D cubes. In Fig. 1 we show the T-$\rho$ (top) and T-$\tau$ (bottom) structures, where $\tau$ is the optical depth, of the MS (left, panels A and C) and of the sub-giant (right, panels B and D). As a representative case, we show both models at $\rm[Fe/H]=-5$ to investigate their properties and to illustrate the typical behaviour observed across the full set of models. All panels illustrate the respective atmosphere between 4-4 and 0 in $\rm\log\tau_{500}$.
@@
Comparing the 3D structures with their 1D counterparts, the dominant differences are apparent towards the upper layers of the atmospheres. For the MS model (Fig. 1, panels A and C), the mean 3D temperature structure starts to deviate from that of the 1D HE model around optical depth $\rm\log\tau_{500}\sim-2$. The $\rm 3D-1D$ difference reaches values of $\rm\sim-1\,500\ K$ at $\rm\log\tau_{500}=-4$. In those layers, the 3D RHD model is significantly and systematically cooler than the 1D MARCS model, which is well known and was extensively discussed in the earlier work on 3D models of very metal-poor MS and red giant stars (Gallagher et al. 2016, 2017; Collet et al. 2018). This difference is due to the fundamental property of convection, a process that is inherent to 3D RHD models, but is missing or parameterised by the mixing-length theory (Gustafsson et al. 2008) in 1D HE models. In metal-poor stars, the atmospheric structure is dominated by adiabatic cooling of rising, and hence expanding, fluid parcels of the gas. This is in stark contrast to stars with solar-like metallicity: in such conditions, the high metallicity leads to significant radiative heating (Stein & Nordlund 1998), which counteracts the adiabatic cooling effect and implies that the average 3D structure remains nearly in radiative equilibrium (e.g. Nordlund et al. 2009; Freytag et al. 2012; Magic et al. 2013a, b).
@@
A very similar behaviour can be seen in the structure of the 3D model of the sub-giant star (Fig. 1, panels B and D). The 3D RHD model is over $\rm\sim-1\,000\ K$ cooler compared to its MARCS counterpart in the layers with $\log\tau_{500}\lesssim-3$. Although the effect is very large, it is slightly less than that of the MS 3D model in the uppermost layers ($\log\tau_{500}\sim-4$). One would naively expect that a sub-giant 3D model, due to lower pressures, would show larger deviations from the 1D HE model compared to the MS case. However, there is another interesting effect. The lower density and cooler structure of this model favours formation of different molecules, including CH, OH, and CO. Due to radiative re-heating, the temperature balance of the 3D model hence is slightly closer to that of 1D HE MARCS model. We emphasise that this difference is not a direct consequence of the C-enhancement, it is also present in models with scaled-solar composition, but of the overall absence of metal absorbers in the energy budget of the atmosphere.
3.1.2 3D CEMP versus 3D scaled-solar
@@
Comparing C-enhanced and scaled-solar 3D models directly (Fig. 1, blue crosses and solid blue lines, respectively), we find that for the very metal-poor MS star (panels A and C), the effect is modest compared to the $\rm 3D-1D$ differences described above. At optical depth of $\log\tau_{500}>-3$, there is no significant difference, while towards the upper boundary of the atmosphere, the mean temperature of the CEMP model is $\rm\sim 400\ K$ hotter at $\rm\log\tau_{500}\sim-4$. In the sub-giant model (panels B and D), the same effect of C-enhancement is noticeable already at $\rm\log\tau_{500}\sim-2$, with temperature differences of $\rm\sim 200-300\ K$ throughout the entire optically thin part of the atmosphere. Overall, the effect on the MS model is $\rm\sim 100\ K$ larger at around $\rm log\ \tau_{500}\lesssim-4$ than on the sub-giant. Towards deeper layers, however, $\rm\log\tau_{500}\gtrsim-3$, the scaled-solar sub-giant remains consistently cooler than its CEMP counterpart, while the departures between the MS models are already negligible.
@@
The reason for this additional heating in the CEMP models is found in the increased abundance of C, which implies increased CH, CO, and CN molecular abundances, a higher molecular line opacity, and consequently more radiative heating of the upper atmospheric layers. Because O – as well as N – remain at their scaled-solar values, specifically the CH G band is one of the main contributors to the line opacity beyond $\rm\log\tau_{500}\sim-2$. In cooler atmospheres with very low [Fe/H], this effect is most noticeable. First of all, this is because already in the scaled-solar approximation C is an order of magnitude more abundant than Fe. In addition, the CH number densities are higher in cooler systems, as the chemical equilibrium favours a higher $\rm CH/C$ for lower T values. The re-heating of the upper atmospheric layers induced by the C-enhancement hence produces the opposite of the $\rm 3D-1D$ effect described in Sect. 3.1.1. We furthermore note that there is no such effect in 1D HE MARCS models. To highlight this, we include one example CEMP MARCS model with the same stellar parameters in panels A and C.
@@
Because of its lower $\rm log(g)$, the sub-giant model overall has a lower density at the same temperature as the MS model. A consequence of this is that the formation height of spectral lines generally shifts towards deeper layers in terms of optical depth. This explains the larger departures ($\rm 3D\ CEMP-\text{scaled-solar}$ models) at optical depth $\rm\log\tau_{500}\gtrsim-3$ among the sub-giant models compared to the MS models (compare the difference between blue lines and blue crosses in panel C with panel D) through the shift of the molecular line heating contribution to deeper atmospheric layers. Generally, decreasing the overall effective temperature causes the opposite effect, as the increased molecular number densities increase the opacity and hence shift the contribution function to lower optical depth (but see also the discussion in Appendix A).
@@
Figure 3: Difference between 3D CEMP and 1D MARCS models as a function of metallicity. Top: Relative temperature difference (in percent). Middle: Logarithmic density differences (in dex). Bottom: CH/H number density fraction, where $\rm n^{*}=n_{CH}/n_{H}$ is the CH number density relative to the hydrogen number density. Dashed blue lines correspond to $\rm T_{eff}=5750\ K,\ log(g)=4.5$ models, solid red curves to $\rm T_{eff}=5250\ K,\ log(g)=3.0$. The differences in each figure are computed at the representative formation height of the strong lines in the G band. See Sect. 3.2 for more information.
@@
Figure 4: CH G band computed with M3DIS 3D and 1D MARCS model atmospheres. Solid black lines represent the CEMP 3D model, dashed blue lines the corresponding scaled-solar 3D model. Solid red lines show the 1D MARCS model with the same stellar parameters. Top: $\rm T_{eff}=5750\ K,\ log(g)=4.5$. Bottom: $\rm T_{eff}=5250\ K,\ log(g)=3.0$.
3.1.3 3D CEMP versus 1D HE
@@
Combining the results from Sects. 3.1.1 and 3.1.2, we next compared our M3DIS 3D CEMP model atmospheres with 1D HE scaled-solar MARCS models. The difference between these two model types is particularly important, as many studies rely on 1D scaled-solar MARCS models for spectral synthesis and abundance analysis. This raises the question of whether this simplifying assumption introduces systematic effects in the derived abundances.
@@
To find the answer, we first show the full 3D temperature structure of both selected 3D CEMP models at $\rm[Fe/H]=-5$ in Fig. 2. For both models, MS (left-hand side) and sub-giant (right-hand side), we used the C-enhancement of $\rm[C/Fe]=+3$. We furthermore show in the inset a horizontal slice of the same atmospheres, interpolated to the optical surface at $\rm\log\tau_{500}=0$. We note the very different scales and granular size between the MS (left panel) and the sub-giant (right panel) models. The model of the MS star is, because of a much stronger gravity, about 20 times smaller than that of a sub-giant simulation, the latter exceeding 100 Mm in size. As described in the previous sections, the difference between 3D CEMP and 1D MARCS models is most striking in the optically thin layers of the atmosphere, where 3D CEMP models are still significantly cooler, even though the difference is reduced by the C-enhancement-induced re-heating. The two models converge towards the optical surface; the agreement is best in the proximity of $\rm\tau=2/3$, indicating a very similar effective temperature of 1D and 3D models.
@@
Towards the bottom of Fig. 2 we furthermore include horizontal bars, indicating the formation region of the G band. For this, we selected a representative CH line, compute the contribution function in its core for each model atmosphere, and used the location of its peak as proxy for the optical depth at which the G band typically forms. We drew the horizontal lines until $\rm\tau_{500}=1$, which corresponds to the formation of weaker features, line wings and the continuum. We emphasise that the computation of the G-band formation region does not affect the results or conclusions of this study; it is intended solely as a useful reference to aid in the analysis of model differences and to support their physical interpretation. In Fig. 2 we see that the G band forms slightly deeper within the atmosphere of the sub-giant as compared to the MS model, which is a consequence of the significantly lower surface gravity and hence overall lower density. This shift happens despite the lower effective temperature of the 3D sub-giant model. In the 1D model the upper layers are strongly dominated by radiative equilibrium – and hence the effective temperature – which means the drop in temperature between the MS and sub-giant models is more significant, the opacity rises and the formation region of the G band shifts in the opposite direction.
@@
The evolution of the differences between 3D CEMP and 1D MARCS models with metallicity can be investigated in Fig. 3. For clarity, we refrain from showing the differences between the full profiles as a function of optical depth for each metallicity. Instead, we present a representative 3D–1D difference for each metallicity, computed from the atmospheric regions most relevant to G-band formation. Accordingly, the data shown in this figure are evaluated at the G-band formation height in each respective atmosphere, as described in the previous paragraph.
@@
In panel A of Fig. 3 the relative temperature difference is shown for the sub-giant models (solid red lines) and for the MS models (dashed blue lines). We find that 3D CEMP models are on average between $20\%$ and $25\%$ cooler than 1D HE models. At metallicities 6-6, 5-5, and 4-4, there is no significant variation in the 3D–1D difference. At $\rm[Fe/H]=-3$ the differences start to decline, and reach values around $5\%$ at metallicity 2-2. Besides forming in cooler parts of the atmosphere, the G band furthermore forms in regions of lower densities (panel B), with $\rm\log\rho_{3D}/\rho_{1D}$ between 0.6-0.6 dex at $\rm[Fe/H]=-6$ and 0.3-0.3 dex at $\rm[Fe/H]=-2$. Note that this is true despite the fact that metal-poor 1D HE models generally are less dense at the same optical depth (due to the hotter temperatures) and the data in panel B mostly reflect the differences in formation height through the exponential growth of density with depth.
@@
Additionally, the MS models show 3D CEMP effects similar to – if not even slightly larger than – the sub-giant models, especially when considering the temperature differences (Fig. 3, panel A). However, at this stage it is important to consider that the molecular equilibrium, which ultimately determines the CH number densities and hence directly influences the strength of the spectral lines in the G band, is not a linear function of temperature and density, but rather exhibits a temperature and density dependence as dictated by the Saha equation (e.g. Carson 1992):
@@
where $\rm n_{x}$ is the number density of species $\rm x$, $\rm z_{x}$ the partition functions and $\rm D_{CH}$ the dissociation energy of the CH molecule, which is set to $\rm 3.47\ eV$ (Popa et al. 2023; Storm et al. 2025, and references therein). To highlight this non-linear dependence, we include panel C, which shows the fraction of CH/H number densities $\rm n^{*}_{3D}/n^{*}_{1D}$ – where $\rm n^{*}=n_{CH}/n_{H}$ – in the same formation region as panels A and B. Here it becomes obvious that the number density of CH/H molecules in the sub-giants deviate by a factor of $\rm\sim 1.2$ more between 3D and 1D than what is found in MS models. So despite showing a slightly smaller 3D effect in temperature and density, the 3D effect is more prominent in the CH/H number density of sub-giants, which will ultimately be reflected in the G band and $\rm[C/Fe]$ abundance corrections (see Sect. 4).
@@
Overall, with increasing [Fe/H], the differences in the structure between 3D CEMP and 1D HE models decrease noticeably. The reason is that with increasing metal content, radiative heating starts to dominate the overall energy budget of the photosphere. The consequence is that adiabatic cooling becomes less dominant in the outer layers, which ultimately closes the gap between the average 3D structure and 1D HE MARCS models, which do not include this cooling mechanism in the first place.
@@
Our results are in a good agreement with work presented by Gallagher et al. (2016, 2017), although a different code (CO5BOLD: Freytag et al. 2012) was used in that study and their choice of stellar parameters does not fully overlap with that of our models. Gallagher et al. (2017) furthermore note that if the O abundance is changed, this has an impact on the number densities of the CO molecule. Consequently, the amount of free C available to form CH is decreased. Whereas generally it would be of interest to test non-solar abundances of different elements, in this work we did not explore the dependence on the $\rm C/O$ ratio. First, because there is evidence of CEMP stars with only very modest O enhancement (Bessell et al. 2004; Iwamoto et al. 2005). Also, most diagnostic OH features that are suitable for O abundance measurements (Asplund & García Pérez 2001) are located in the UV range, which is extremely sensitive to non-LTE (NLTE) effects in metal-poor stars (Bergemann & Nordlander 2014; Lind & Amarsi 2024). Furthermore, no 3D NLTE $\rm C/O$ ratios are available for CEMP stars (Collet et al. 2018). We also note that Gallagher et al. (2017) focus on a hot sub-giant with metallicity 3-3, which leads to smaller differences between 1D and 3D models – partly because of the higher metallicity, and partly because of the lower $\rm CH/C$ ratio in both CEMP and non-CEMP models. We explore the impact of different metallicities and stellar parameters on synthetic stellar spectra and C abundance determinations in the following sections.
3.2 Spectrum synthesis of the CH G band